OpenStudy (anonymous):

solve y=2x-5, y=6x+6 using systems of equations. Leave all answers in fractional form.

OpenStudy (anonymous):

if y equals both 2x-5 AND 6x+6, what does that tell us?

OpenStudy (anonymous):

um.. i dont know....

OpenStudy (anonymous):

think victoria. 3 equals both 6/2 and 18/6 what does this tell us about the relationship between 6/2 and 18/6?

OpenStudy (anonymous):

That they are common factors? Im doing summer homework and i dont remember any of this...

OpenStudy (anonymous):

no, don't think of anything you learned at school about math. if 3 = 6/2 and 3 = 18/6, this tells us that 6/2 = 18/6 since they are both equal to 3. understand?

OpenStudy (anonymous):

Yes. But how does that help me with this problem? If i wanted to subsitute it, i would need to move one y so its in the equation, right? then but one of the equations in the other equation?

OpenStudy (anonymous):

it helps you because y equals both 2x-5 and 6x+6 so that must mean that 2x-5 = 6x+6 solve for x and then substitute in one of the original equations to get y

OpenStudy (anonymous):

how do i sovle for x if there is one on either side of the equation?

OpenStudy (anonymous):

okay, let me give you an example. 3x-2 = 4x+4 subtract 3x on both sides 3x-3x-2 = 4x-3x+4 which gives -2 = x +4 subtract 4 on both sides -2 -4 = x+4-4 which gives -6 = x that is, x = -6 do your problem in a similar fashion.

OpenStudy (anonymous):

okay. so i should take 2x-5 = 6x+6 and take 2x off bith sides. So it would be -5=4x+6, then subtract... 6. right?

OpenStudy (anonymous):

that is correct. very good!

OpenStudy (anonymous):

So then it would be -11=4x. so then do i divide -11 by 4?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

but.. i get a decimal? then i put that in one of origonal problems, or both of them?

OpenStudy (anonymous):

read what your question says. leave all answers in fractional form. so leave x = -11/4 and substitute thaT fraction in the equation for y

OpenStudy (anonymous):

um.. im horrible with fractions. would it be y=2(-11/4)-5... so it would be.. y=-22/8 -5?

OpenStudy (anonymous):

\[2 \times \frac{11}{4} = \frac{11}{2}\]

OpenStudy (anonymous):

22/8 is the same as 11/4

OpenStudy (anonymous):

but then how do i subtract the 5?

OpenStudy (anonymous):

y = (-11/2)-5 = -(11/2)-(10/2) because 5 is the same as 10/2 y = -21/2 I suggest you see all those videos I posted if you want to learn this stuff. they will help you learn fractions properly.

OpenStudy (anonymous):

Thanks. Alot. (:

OpenStudy (anonymous):

no problem